OpenStudy (itsbribro):

The sum of two consecutive even integers is -46. what are the numbers? A.) -26 and -20 B.) -22 and -20 C.) -26 and -24 D.) -24 and -22

OpenStudy (nuke_ur_friend):

a

OpenStudy (nuke_ur_friend):

no

OpenStudy (rushwr):

D

OpenStudy (nuke_ur_friend):

wait

OpenStudy (rushwr):

Since it says consecutive even number it can't be A so it should be D

OpenStudy (nuke_ur_friend):

ur right

OpenStudy (rushwr):

\[-22 - 24 = -46\]

OpenStudy (itsbribro):

i was gonna say D or A lol but i was confused

OpenStudy (campbell_st):

well let the 1st integer even be x then the next consecutive even is x + 2 so then the sum is x + x + 1 = -46 or 2x + 2 = -46 now solve for x

OpenStudy (rushwr):

Thanks @Nuke_ur_friend

OpenStudy (itsbribro):

thanks @Rushwr & @campbell_st